# Vortex Avalanches and Collective Motion in Neutron Stars [^1] Released on March, 1st, 2024

### v3

@@ 1 Introduction
The simplest superfluid model that self-consistently describes vortex pinning is the Gross–Pitaevskii (GP) model, and this has previously been used to study rotational glitches in the neutron star crust (Warszawski & Melatos, 2011; Warszawski et al., 2012; Melatos et al., 2015; Lönnborn et al., 2019; Poli et al., 2023). Given the huge disparity between the length scales of individual vortex cores ($\sim 100\,$fm) and the star itself ($\sim 10$ km), such models can only resolve a much smaller system, and previous studies have been limited to $\lesssim 200$ vortices. Although these models have reproduced some features of pulsar glitches, it is unclear whether the results can be scaled up to the true parameter regime. Indeed, Warszawski & Melatos (2011) found that glitch size decreased as they increased the number of vortices, and so their results focussed on simulations with $<100$ vortices.
An alternative approach is the point-vortex (or vortex-filament) model (Howitt et al., 2020; Cheunchitra et al., 2024), which tracks only the position of vortices rather than the density and velocity of the superfluid. This is computationally more efficient and such models have produced glitches in systems of up to 5,000 vortices. However, the interactions between vortices, and their interactions with pinning sites, cannot be treated fully in such a model, and must instead be parameterised. Moreover, the dynamics depend qualitatively on the degree of dissipation in the model, which must be incorporated in an ad hoc manner, and greatly exceeds the dissipation expected in a neutron star.
We study the dynamics of up to $600$ vortices in a two-dimensional GP model, in the presence of a spinning down crust. We show that, if the spin down is sufficiently slow, then rotational glitches do occur and are associated with vortex avalanches. We also demonstrate that the results are essentially independent of the degree of dissipation, provided that this is made sufficiently small.
@@ 2.1 Numerical Model
where $m=2m_{\mathrm{n}}$ is the mass of a superconducting particle (i.e. a neutron Cooper pair), $V(\mathbf{r},t)$ is an imposed potential, $g$ measures the self-repulsion of the superfluid, $\boldsymbol{\Omega}(t)$ is the angular velocity of the reference frame, and $\hat{\mathbf{L}}$ is the angular momentum operator. We will adopt Cartesian coordinates $(x,y,z)$ such that $\boldsymbol{\Omega}=\bigl{(}0,0,\Omega(t)\bigr{)}$, and so $\boldsymbol{\Omega}\cdot\hat{\mathbf{L}}=-\mathrm{i}\hbar\,\Omega\,\partial/\partial\phi$, where $\phi$ is the angular coordinate around the $z$-axis. We solve Eq. (1) numerically in the $xy$-plane, assuming that $\psi$ has no $z$-dependence; for details see Sec. 2.2.
@@
$\int\mathrm{d}^{2}\mathbf{r}\,\psi^{\ast}\left(\hat{H}_{\mathrm{GP}}-\mu\right)\psi,$ (7)
decreases monotonically with time. In the context of ultracold quantum gases, $\gamma$ arises from the interaction between superfluid and normal components at nonzero temperatures, and typically has a value $\lesssim 10^{-3}$ (e.g. Bradley et al., 2008; Blakie et al., 2008; Rooney et al., 2012). In the inner crust of a neutron star, the true dissipation mechanisms are more complicated, and $\gamma$ serves only as a crude parameterisation. Nonetheless, previous studies have typically adopted values $\gamma\geqslant 0.02$ (Warszawski & Melatos, 2011; Warszawski et al., 2012; Melatos et al., 2015; Lönnborn et al., 2019). We aim to determine how small $\gamma$ must be such that it plays little (if any) role in the qualitative behaviour of the system.
@@ 2.2 Numerical Procedure
Simulations are prepared by first evolving in imaginary time, i.e. by setting $\gamma=0$, $\Omega=\Omega_{0}$, and replacing $t\rightarrow\mathrm{i}t$ in Eq. (1(e.g. Modugno et al., 2003), beginning with a random phase at each grid point. After evolving in imaginary time for a sufficiently long period (typically $>$7,500 $\tau$), the wavefunction achieves a quasi-equilibrium where the superfluid density and angular momentum are essentially steady. This is taken as the initial condition for the damped GP equation.
Each vortex is associated with a quantum of circulation $2\pi\hbar/m$. In the absence of pinning sites we would therefore expect the initial number of vortices to be roughly $N_{v}\simeq\Omega_{0}R_{\mathrm{con}}^{2}m/\hbar=(\Omega_{0}\tau)(R_{\mathrm{con}}/\xi)^{2}$, so that the average rotation rate of the superfluid is roughly $\Omega_{0}$. For example, with $R_{\mathrm{con}}=230\xi$ and $\Omega_{0}=2\pi\times 10^{-3}\tau^{-1}$ we would expect $N_{v}\simeq 332$. To study significantly more vortices than previous GP studies we take $\Omega_{0}=2\pi\times 10^{-3}\tau^{-1}$ or $\Omega_{0}=4\pi\times 10^{-3}\tau^{-1}$. In practice, the presence of pinning sites means that there are many different quasi-equilibrium states that can be achieved during imaginary time propagation, and the actual number of vortices can differ from this prediction by up to 20%. Even though the number of pinning sites greatly exceeds the number of vortices, there are usually a small number ($\lesssim 5\%$) of vortices that are unpinned in the initial state (see the bottom-left inset of Fig. 1 (a), for example). Following (Liu et al., 2024), we define a vortex to be pinned if it is located within $1.25w$ of a pinning site.
Adopting a characteristic value of $\xi=200$ fm for the coherence length in the crust (Graber et al., 2017), we find that $\mu\equiv\hbar^{2}/m\xi^{2}\simeq 518\,\text{eV}$ and $\tau\equiv\hbar/\mu\simeq 1.3\times 10^{-18}$ s. For computational feasibility, the domain size and angular velocity used are far from the typical values for a neutron star. A comparison between the estimated physical parameters for real neutron stars and the parameters used in our simulations is presented in Appendix A. Our objective in the present work is to include as many vortices as feasible to study any resulting collective dynamics, which requires compromise with regard to the other physical parameters.
@@ 3.1 Macroscopic Observables
Figure 1: Spin-down dynamics for $\Omega_{0}=4\pi\times 10^{-3}\tau^{-1}$, $\dot{\Omega}=2.5\pi\times 10^{-8}\tau^{-2}$ and $\gamma=5\times 10^{-3}$ over a time span of $4.45\times 10^{4}\tau$. (a) Time series of $N_{v}$ (black solid line), $\langle L_{z}\rangle$ (green dashed line), and $\langle L_{z}\rangle_{v}$ (green dotted line). Glitches, i.e. sudden jumps in $\langle L_{z}\rangle$, are labelled as G.1 to G.9 and color coded by time. The bottom-left inset shows the initial vortex locations, whether pinned (blue filled circles) or unpinned (red hollow circles). The pinning sites are at the vertices of the square grid shown. The top-right inset shows a close-up of the first glitch (G.1). (b) Vortex trajectories in the rotating frame, colour-coded by time. The black dashed circle indicates the region $r\leqslant 210\xi$ within which analysis is performed, i.e. we take $R=210\xi$ in Eqs. (8) and (9).
@@
To count the number of vortices we first determine their locations, at a given time, by interpolating the superfluid velocity to a sub-grid scale and identifying points of singularity. Identifying vortices is problematic close to the edge of the container, $r=R_{\mathrm{con}}$, where the density vanishes and “ghost vortices” often arise. Hence, our subsequent analysis is performed within the subdomain $|\mathbf{r}|\leqslant R=210\xi\simeq 0.91R_{\mathrm{con}}$.
@@
The motion of vortices is illustrated in Fig. 1 (b), which plots vortex trajectories (in the spinning down frame) color coded according to time as in Fig. 1 (a). Each glitch is associated with the unpinning and outward migration of multiple vortices, localized in both time and space, which we interpret as a vortex avalanche. Each avalanche occurs within a narrow channel that is aligned roughly in the radial direction. However, individual vortex trajectories are not purely radial, and most follow roughly circular arcs in a clockwise direction, as shown in Fig. 2. This glitching behaviour occurs only if the spin-down rate, $\dot{\Omega}$, and dissipation, $\gamma$, are sufficiently small (see Appendix B).
@@ 3.2 Vortex Avalanches
In practice, the depinning of a vortex in the GP model is more complicated than in this simplified description, and the critical value can vary by about $20\%$ (e.g. Liu et al., 2024). For the values of $V_{0}$ and $w$ used here, the critical velocity is $v_{c}\simeq 0.2\xi/\tau$, this can help anticipate depinning events. Fig. 2 presents snapshots of the vortex locations in the vicinity of the first glitch (G.1) from Fig. 1. Vortices are coloured according to $|\vec{v}_{s,j}|$ ( Eq. (11)), where the sum is taken over all vortices in the system. By $t=\text{8,500}\,\tau$ a few vortices have reached the threshold, $|\vec{v}_{s,j}|>v_{c}$, at which depinning is expected; the first vortex to depin is labelled with a black arrow. This vortex has a “knock-on” collision with another pinned vortex, causing it to depin (see supplementary movie). As vortices migrate outward, the residual Magnus force on other pinned vortices increases, as shown in Fig. 2(a)(iii), causing further depinning.
@@
where $W$ is a smoothing kernel (Monaghan, 1992) of width $h$. We choose the value of $h$ based on the average distance between vortices, with $h(t)=2.45\sqrt{\pi R_{\mathrm{con}}^{2}/N_{v}(t)}$. Figure  2(b) shows the subsequent dynamics, including the post-glitch period. As a result of vortices exiting the domain, a void appears in the (coarsened) vorticity. Multiple vortices depin and orbit clockwise around this void, causing the void to propagate inward and gradually disperse. By $t=\text{13,000}\,\tau$, as shown in Fig. 2(b)(iv), all vortices have Magnus forces below the threshold for depinning. We observe similar dynamics in each of the subsequent glitches.
The vortex trajectories shown in Fig. 2 suggest that vortices behave collectively both during and after the glitch. In Appendix C, we confirm this by introducing an order parameter that quantifies the local correlation of the vortex motions.
4 Conclusions
Simulation Neutron Star Neutron Star
(dimensionless) (dimensional) (dimensionless)
$R_{\mathrm{con}}$ 230 $10$ km $10^{13}$
$w$ 1 10 fm $0.05$
$V_{0}$ 2 $1$ MeV $2\times 10^{3}$
$d_{p}$ 10 $10$ fm $0.05$
$\Omega_{0}$ $(2,4)\pi\times 10^{-3}$ $(10^{0},10^{3})$ s-1 $(10^{-18},10^{-15})$
$\dot{\Omega}$ $2.5\pi\times(10^{-8},10^{-7})$ $(10^{-24},10^{-5})$ s-2 $(10^{-60},10^{-41})$
Appendix A Numerical vs. Neutron Star Parameters
In the GP model, the characteristic scales of length, time and energy are related. Taking the coherence length in the crust as a guide, $\xi=200$ fm  (Graber et al., 2017), one can find that the characteristic energy scale is
$\mu\equiv\frac{\hbar^{2}}{m\xi^{2}}\simeq 518\,\text{eV}$ (A1)
and
$\tau\equiv\frac{\hbar}{\mu}\simeq 1.3\times 10^{-18}\,s.$ (A2)
Then one can compare the relevant parameters of neutron stars in the dimensionless scale in Table 1.
Appendix B Parameter dependence in GP modeling
The results of Sec. 3 demonstrate the same kind of vortex avalanche behavior believed to occur in neutron stars. However, given that it is not possible to replicate the true parameter conditions of a neutron star in the computational model, it is important to determine the extent to which the results depend on the key parameters: the initial rotation rate, $\Omega_{0}$, the spin-down rate, $\dot{\Omega}$, and the dissipation, $\gamma$.
@@ Appendix B Parameter dependence in GP modeling
Finally, we investigate the role of dissipation by decreasing or increasing $\gamma$ by a factor of 10. In the simulation with dissipation decreased to $\gamma=5\times 10^{-4}$ we observe similar glitching behavior, and the typical magnitude of the glitches is not significantly affected. By contrast, in the simulation with dissipation increased to $\gamma=5\times 10^{-2}$ we do not observe glitches. We conclude from this that glitches can only occur when the level of dissipation is sufficiently small, and that once it is sufficiently small it does not play a significant role in the glitch dynamics.
Appendix C Vortex Collective Motion
In order to quantify whether the vortices are exhibiting collective motion, we introduce an order parameter,
$\varphi(r,t)=\displaystyle\frac{1}{N_{v,\mathrm{moving}}}\sum_{j=1}^{N_{v,\mathrm{moving}}}\left|\vec{u}_{j}\right|,$ (C1)
$\vec{u}_{j}(r,t)=\frac{1}{\mathcal{N}_{j}}\sum_{k\in C_{j}}\frac{\Delta\vec{r}_{j}/\Delta t}{|\Delta\vec{r}_{j}/\Delta t|},$ (C2)
where $C_{j}=\{k\neq j:|\vec{r}_{j}(t)-\vec{r}_{k}(t)|\leqslant r\}$, and $\mathcal{N}_{j}$ is the number of vortices in $C_{j}$. The locations of vortices are tracked every one unit of time, namely, $\Delta t=\tau$, allowing us to identify the vortex motion with sufficiently fine time resolution.
Fig. 4 presents a time series of $\varphi$ for the same simulation shown in Fig. 1. As in Fig. 1, the start and end times of each glitch are indicated by vertical dashed and dotted lines, respectively. We see that periods of collective motion, indicated by values of $\varphi(r,t)\gtrsim 0.5$ over a range of $r$, typically occur throughout most of the glitch, and often continue significantly after the glitch.

### v4

@@ 1 Introduction
The simplest superfluid model that self-consistently describes vortex pinning is the Gross–Pitaevskii (GP) model, and this has previously been used to study rotational glitches in the neutron star crust (Warszawski & Melatos, 2011; Warszawski et al., 2012; Melatos et al., 2015; Lönnborn et al., 2019; Verma et al., 2022; Poli et al., 2023; Shukla et al., 2024). Given the huge disparity between the length scales of individual vortex cores ($\sim 10\,$fm) and the star itself ($\sim 10$ km), such models can only resolve a much smaller system, and previous studies have been limited to $\lesssim 200$ vortices. Although these models have reproduced some features of pulsar glitches, it is unclear whether the results can be scaled up to the true parameter regime. Indeed, Warszawski & Melatos (2011) found that glitch size decreased as they increased the number of vortices, and so their results focussed on simulations with $<100$ vortices.
An alternative approach is the point-vortex (or vortex-filament) model (Howitt et al., 2020; Cheunchitra et al., 2024), which tracks only the position of vortices rather than the density and velocity of the superfluid. This is more computationally efficient and such models have produced glitches in systems of up to 5,000 vortices. However, the interactions between vortices, and their interactions with pinning sites, cannot be treated fully in such a model, and must instead be parameterised. Moreover, the dynamics depend qualitatively on the degree of dissipation in the model, which must be incorporated in an ad hoc manner, and greatly exceeds the dissipation expected in a neutron star.
We study the dynamics of up to $600$ vortices in a two-dimensional GP model, in the presence of a spinning down crust. We show that, if the spin down is sufficiently slow, then rotational glitches do occur and are associated with vortex avalanches and collective motions.. We also demonstrate that the results are essentially independent of the degree of dissipation, provided that this is made sufficiently small.
@@ 2.1 Numerical Model
where $m=2m_{\mathrm{n}}$ is the mass of a superconducting particle (i.e. a neutron Cooper pair), $V(\mathbf{r},t)$ is an imposed potential, $g$ measures the self-repulsion of the superfluid, $\boldsymbol{\Omega}(t)$ is the angular velocity of the reference frame, and $\hat{\mathbf{L}}$ is the angular momentum operator. We will adopt Cartesian coordinates $(x,y,z)$ such that $\boldsymbol{\Omega}=\bigl{(}0,0,\Omega(t)\bigr{)}$, and so $\boldsymbol{\Omega}\cdot\hat{\mathbf{L}}=-\mathrm{i}\hbar\,\Omega\,\partial/\partial\phi$, where $\phi$ is the angular coordinate around the $z$-axis. We solve Eq. (1) numerically in the $xy$-plane, assuming that $\psi$ has no $z$-dependence; for details see Sec. 2.2. This reduction to two dimensions precludes dynamics that is intrinsically three-dimensional, such as vortex reconnections and Kelvin waves (e.g. Barenghi et al., 2001; Link & Levin, 2022). However, the model still self-consistently incorporates pinning and depinning, which are the essential processes in glitches, and allows hundreds of vortices to be simulated in a computationally feasible timescale. Future studies will be required to quantify the role of three dimensional dynamics in the glitch process.
@@
$\int\mathrm{d}^{2}\mathbf{r}\,\psi^{\ast}\left(\hat{H}_{\mathrm{GP}}-\mu\right)\psi.$ (7)
In the context of ultracold quantum gases, $\gamma$ arises from the interaction between superfluid and normal components at nonzero temperatures, and typically has a value $\lesssim 10^{-3}$ (e.g. Bradley et al., 2008; Blakie et al., 2008; Rooney et al., 2012). In the inner crust of a neutron star, the true dissipation mechanisms are more complicated, and $\gamma$ serves only as a crude parameterisation. Nonetheless, previous studies have typically adopted values $\gamma\geqslant 0.02$ (Warszawski & Melatos, 2011; Warszawski et al., 2012; Melatos et al., 2015; Lönnborn et al., 2019). We aim to determine how small $\gamma$ must be such that it plays little (if any) role in the qualitative behaviour of the system.
Simulation Neutron Star Neutron Star
(dimensionless) (dimensional) (dimensionless)
$R_{\mathrm{con}}$ 230 $10$ km $10^{18}$
$w$ 1 10 fm $1$
$V_{0}$ 2 $1$ MeV $4.8$
$d_{p}$ 10 $10$ fm $1$
$\Omega_{0}$ $(2,4)\pi\times 10^{-3}$ $(10^{0},10^{3})$ s-1 $(10^{-21},10^{-18})$
$\dot{\Omega}$ $2.5\pi\times(10^{-8},10^{-7})$ $(10^{-24},10^{-5})$$s^{-2}$ $(10^{-65},10^{-46})$
@@ 2.2 Numerical Procedure
Simulations are prepared by first evolving in imaginary time, i.e. by setting $\gamma=0$, $\Omega=\Omega_{0}$, and replacing $t\rightarrow\mathrm{i}t$ in Eq. (1(e.g. Modugno et al., 2003), beginning with a random phase at each grid point. After evolving in imaginary time for a sufficiently long period (typically $>$7,500 $\tau$), the wavefunction achieves a quasi-equilibrium in which the superfluid density and angular momentum are essentially steady. This is taken as the initial condition for the damped GP equation.
Each vortex is associated with a quantum of circulation $2\pi\hbar/m$. In the absence of pinning sites we would therefore expect the initial number of vortices to be roughly $N_{v}\simeq\Omega_{0}R_{\mathrm{con}}^{2}m/\hbar=(\Omega_{0}\tau)(R_{\mathrm{con}}/\xi)^{2}$, so that the average rotation rate of the superfluid is roughly $\Omega_{0}$. For example, with $R_{\mathrm{con}}=230\xi$ and $\Omega_{0}=2\pi\times 10^{-3}\tau^{-1}$ we would expect $N_{v}\simeq 332$. To study significantly more vortices than previous GP studies we take $\Omega_{0}=2\pi\times 10^{-3}\tau^{-1}$ or $\Omega_{0}=4\pi\times 10^{-3}\tau^{-1}$. In practice, the presence of pinning sites means that there are many different quasi-equilibrium states that can be achieved during imaginary time propagation, and the actual number of vortices can differ from this prediction by up to 20%. Even though the number of pinning sites greatly exceeds the number of vortices, there are usually a small number ($\lesssim 5\%$) of vortices that are unpinned in the initial state (see the bottom-left inset of Fig. 1 (a), for example). Following Liu et al. (2024), we define a vortex to be pinned if it is located within $1.25w$ of a pinning site.
2.3 Numerical vs. Neutron Star Parameters
Adopting a characteristic value of $\xi=10$ fm for the coherence length in the crust (e.g. Mendell, 1991; Seveso et al., 2016; Graber et al., 2017), we find that
$\mu\equiv\frac{\hbar^{2}}{m\xi^{2}}\simeq 207\,\text{keV}$ (8)
and
$\tau\equiv\frac{\hbar}{\mu}\simeq 3.2\times 10^{-21}\,\textrm{s}.$ (9)
A comparison between the parameter values used in our simulations and those estimated for real neutron stars is given in Table  1. For computational feasibility, the domain size and angular velocity we use are far from the typical values for a neutron star. Our objective in the present work is to include as many vortices as feasible to study any resulting collective dynamics, which requires compromise with regard to the other physical parameters.
@@ 3 Spin-down dynamics
Figure 1: Spin-down dynamics for $\Omega_{0}=4\pi\times 10^{-3}\tau^{-1}$, $\dot{\Omega}=2.5\pi\times 10^{-8}\tau^{-2}$ and $\gamma=5\times 10^{-3}$ over a time span of $4.45\times 10^{4}\tau$. (a) Time series of $N_{v}$ (black solid line), $\langle L_{z}\rangle$ (green dashed line), and $\langle L_{z}\rangle_{v}$ (green dotted line). Glitches, i.e. sudden jumps in $\langle L_{z}\rangle$, are labelled as G.1 to G.9 and color coded by time. The bottom-left inset shows the initial vortex locations, whether pinned (blue filled circles) or unpinned (red hollow circles). The pinning sites are at the vertices of the square grid shown. The top-right inset shows a close-up of the first glitch (G.1). (b) Vortex trajectories in the rotating frame, colour-coded by time. The black dashed circle indicates the region $r\leqslant 210\xi$ within which analysis is performed, i.e. we take $R=210\xi$ in Eqs. (10) and (11). An animation of the same simulation is included in Appendix A.
@@ 3.1 Macroscopic Observables
To count the number of vortices we first determine their locations, at a given time, by interpolating the superfluid velocity to a sub-grid scale and identifying points of singularity. Identifying vortices is problematic close to the edge of the container, $r=R_{\mathrm{con}}$, where the density vanishes and “ghost vortices” often arise. Therefore, our subsequent analysis is performed within the subdomain $|\mathbf{r}|\leqslant R=210\xi\simeq 0.91R_{\mathrm{con}}$.
@@
The motion of vortices is illustrated in Fig. 1 (b), which plots vortex trajectories (in the spinning down frame) color coded according to time as in Fig. 1 (a). Each glitch is associated with the unpinning and outward migration of multiple vortices, localized in both time and space, which we interpret as a vortex avalanche. Each avalanche occurs within a narrow channel that is aligned roughly in the radial direction. However, individual vortex trajectories are not purely radial, and most follow roughly circular arcs in a clockwise direction, as shown in Fig. 2. This glitching behaviour occurs only if the spin-down rate, $\dot{\Omega}$, and dissipation, $\gamma$, are sufficiently small (see Sec. 5).
@@ 3.2 Vortex Avalanches
In practice, the depinning of a vortex in the GP model is more complicated than in this simplified description, and the critical value can vary by about $20\%$ (e.g. Liu et al., 2024). For the values of $V_{0}$ and $w$ used here, the critical velocity is $v_{c}\simeq 0.2\xi/\tau$, this can help anticipate depinning events. Fig. 2 presents snapshots of the vortex locations in the vicinity of the first glitch (G.1) from Fig. 1. Vortices are coloured according to $|\vec{v}_{s,j}|$ ( Eq. (13)), where the sum is taken over all vortices in the system. By $t=\text{8,500}\,\tau$ a few vortices have reached the threshold, $|\vec{v}_{s,j}|>v_{c}$, at which depinning is expected; the first vortex to depin is labelled with a black arrow. This vortex has a “knock-on” collision with another pinned vortex, causing it to depin (see supplementary movie). As vortices migrate outward, the residual Magnus force on other pinned vortices increases, as shown in Fig. 2(a)(iii), causing further depinning. The vast majority of depinning events are attributable to the strength of the Magnus force, with knock-on events being much rarer.
@@
where $W$ is a smoothing kernel (Monaghan, 1992) of width $h$. We choose the value of $h$ based on the average distance between vortices, with $h(t)=2.45\sqrt{\pi R_{\mathrm{con}}^{2}/N_{v}(t)}$. Figure  2(b) shows the subsequent dynamics, including the post-glitch period. As a result of vortices exiting the domain, a void appears in the (coarsened) vorticity. Multiple vortices depin and orbit clockwise around this void, causing the void to propagate inward and gradually disperse. By $t=\text{13,000}\,\tau$, as shown in Fig. 2(b)(iv), all vortices have Magnus forces below the threshold for depinning. We observe similar dynamics in each of the subsequent glitches, which can be seen in the supplementary movie in Appendix A.
4 Vortex Collective Motion
The vortex trajectories shown in Fig. 2 suggest that vortices behave collectively both during and after the glitch. In order to test this, we introduce an order parameter,
$\varphi(r,t)=\displaystyle\frac{1}{N_{v,\mathrm{moving}}}\sum_{j=1}^{N_{v,\mathrm{moving}}}\left|\vec{u}_{j}\right|,$ (15)
$\vec{u}_{j}(r,t)=\frac{1}{\mathcal{N}_{j}}\sum_{k\in C_{j}}\frac{\Delta\vec{r}_{k}/\Delta t}{|\Delta\vec{r}_{k}/\Delta t|},$ (16)
where $C_{j}=\{k\neq j:|\vec{r}_{j}(t)-\vec{r}_{k}(t)|\leqslant r\}$, and $\mathcal{N}_{j}$ is the number of vortices in $C_{j}$. If $\mathcal{N}_{j}<2$ then we omit vortex $j$ from the sum in Eq. (15). The locations of vortices are tracked every one unit of time, namely, $\Delta t=\tau$, allowing us to identify the vortex motion with sufficiently fine time resolution.
Fig. 3 presents a time series of $\varphi$ for the same simulation shown in Fig. 1. As in Fig. 1, the start and end times of each glitch are indicated by vertical dashed and dotted lines, respectively. We see that periods of collective motion, indicated by values of $\varphi(r,t)\gtrsim 0.5$ over a range of $r$, typically occur throughout most of the glitch, and often continue significantly after the glitch. This is consistent with the dynamics illustrated in Fig. 2.
5 Parameter dependence in GP modeling
Our results demonstrate the same kind of vortex avalanche behavior believed to occur in neutron stars. However, given that it is not possible to replicate the true parameter conditions of a neutron star in the computational model, it is important to determine the extent to which the results depend on the key parameters: the initial rotation rate, $\Omega_{0}$, the spin-down rate, $\dot{\Omega}$, and the dissipation, $\gamma$.
@@ 5 Parameter dependence in GP modeling
Finally, we investigate the role of dissipation by decreasing or increasing $\gamma$ by a factor of 10. In the simulation with dissipation decreased to $\gamma=5\times 10^{-4}$ we observe similar glitching behavior, and the typical magnitude of the glitches is not significantly affected. By contrast, in the simulation with dissipation increased to $\gamma=5\times 10^{-2}$ we do not observe glitches. We conclude from this that glitches can only occur when the level of dissipation is sufficiently small, and that once it is sufficiently small it does not play a significant role in the glitch dynamics. In our simulations the critical value of $\gamma$ is comparable to the average rotation rate of the superfluid (measured in units of $\tau^{-1}$), though it likely depends on the rate of spin down, as well as on the choice of pinning potential.
6 Conclusions
Data supporting this publication can be openly accessed under an ‘Open Data Commons Open Database License’ (CC BY 4.0) at https://doi.org/10.25405/data.ncl.25590531.v1.
Appendix A Supplementary Movie
Figure 5: Animation of the same simulation shown in Figs. 13. The color scheme matches that of Fig. 2, except that free vortices are marked by hollow black circles. When the Magnus forces on pinned vortices reaches a threshold, they depin and trigger further depinning, resulting in vortex avalanches. Preview can be seen at https://www.youtube.com/watch?v=dvtMr7haHQo.